When a sound wave or wavelength A. is propagating in a medium, the 'maximum velocity of the particle is equal to the wave velocity. The amplitude of wave is___?
 

Grade:12

1 Answers

Vikas TU
7 years ago
 
When a sound wave or wavelength A. is propagating in a medium, the 'maximum velocity of the particle is equal to the wave velocity. The amplitude of wave cAN BE found as:

Vmax = for particle = Amplitude x w (angular velocity) 
         = A’w

Vwave = w/k = w x A/2pie
Vmax = Vwave
wA/2pie = A'w
A’= A/2pie is the answer.